OpenStudy (anonymous):

What is the slope of the line that passes through the pair of points? (–5.5, 6.1), (–2.5, 3.1) –1 1 –3 3

OpenStudy (anonymous):

point slope method: (3.1 - 6.1) = m(-2.5+5.5) -3=m(3) m=-1 answer is a http://www.freemathhelp.com/point-slope.html